OpenStudy (anonymous):

A uniform rod 0A of length l, resting on smooth surface is slightly disturbed from its vertical position. P is a point on the rod whose locus is a circle during the subsequent motion of the rod. Then the distance OP is equal to.. ?

OpenStudy (anonymous):

plz clear yr qstn by drawing figr

OpenStudy (anonymous):

OP is the radius of the circle

OpenStudy (anonymous):

whats is it equal to in terms of l?

OpenStudy (anonymous):

i think yr qstn is incomplete u have not given locus point co-ordinate........my thinking!!!

OpenStudy (anonymous):

something is missing. OP= l-AP but that's about all i can say and that it is the radius of the circle :x

OpenStudy (anonymous):

i got till that.. but it is solvable.. dont noe how..

OpenStudy (anonymous):

i tried doin with the circle equation and taking the rod as y axis..

OpenStudy (anonymous):

The question is complete and solvable. It was asked in one of our tests. Don't know the answer though, sorry.

OpenStudy (anonymous):

OP= L-AP... You have to state the unknowns with reference to the knowns...even if you can't quantify what they are. the length is known - L- OA is unown (it is L...L=OA) OP is both unknown and the focus of the equation that leaves AP which - whilst remaining unquantified - is the difference between OA (unknown) and OP (unknown)....you could also state for clarity OA-OP= AP....that way you are covering all bases!